OpenStudy (anonymous):

Solve: 1.) 2 log 2 x - log 2 (x+3) = 2 2.)Solve: log 4 x = 1/2 3.)log b 9 = 2 4.)log 2 y = 1/3 (log 2 27)

OpenStudy (anonymous):

1) log2x^2-log(2x+6)=2 log((2x^2)/(2x+6))=2 therfore 10^2=((2x^2)/2x+6)) 100(2x+6)=2x^2 200x+600=2x^2 2x^2-200x-600=0 x^2-100x-300=0 then simplify using the quadratic equation . which by my calculations should give x=102.9150262 or x=-2.915026221